OpenStudy (anonymous):

Please help me! A gas station charges $2.19of gas. Use function notation to describe the relationship between the total cost C(g) and the number of gallons purchased g. (A)C(g)=-2.19g (B)g=2.19C(g) (C)C=g+2.19 (D)C(g)=2.19g

OpenStudy (dan815):

i think it is supposed to say 2.19/g 2.19 per gallon of gas

OpenStudy (dan815):

for every gallon its 2.19 so, the total cost must be the number of gallons * 2.19 Cost = 2.19 * g C(g)=2.19*g <--------------- the C(g) is just called function notation it means cost which is a function of g, telling you that g is a variable

OpenStudy (anonymous):

Yeah sorry its supposed to say $2.19 per gallon! @dan815

OpenStudy (dan815):

okay, so does this make sense now?

OpenStudy (anonymous):

Yes thank you so much! @dan815